### v1

# Long-Tailed Out-of-Distribution Detection with Refined Separate Class Learning

Abstract
Out-of-distribution (OOD) detection is crucial for deploying robust machine learning models. However, when training data follows a long-tailed distribution, the model’s ability to accurately detect OOD samples is significantly compromised, due to the confusion between OOD samples and head/tail classes. To distinguish OOD samples from both head and tail classes, the separate class learning (SCL) approach has emerged as a promising solution, which separately conducts head-specific and tail-specific class learning. To this end, we examine the limitations of existing works of SCL and reveal that the OOD detection performance is sensitive to static temperature value and subject to the impact of informative outliers. To mitigate these limitations, we propose a novel approach termed Refined Separate Class Learning (RSCL), which leverages dynamic class-wise temperature adjustment to modulate the temperature parameter for each in-distribution class and utilizes informative outlier mining to identify diverse types of outliers based on their affinity with head and tail classes. Extensive experiments demonstrate that RSCL achieves superior OOD detection performance while improving the classification accuracy on in-distribution data.
1 Introduction
Deep neural networks (DNNs) are widely known for generating overconfidence predictions when encountering unfamiliar data, leading to the misclassification of out-of-distribution (OOD) samples from unknown classes as members of known classes with high confidence [32, 47]. Such behavior can yield detrimental consequences in safety-critical applications such as autonomous driving, medical diagnosis, and fraud detection [21, 16, 7]. Hence, the task of OOD detection, which focuses on accurately identifying OOD test data while maintaining high performance on in-distribution (ID) data, has received significant attention.
While existing state-of-the-art OOD detection methods [11, 23, 20, 24, 9] have demonstrated remarkable success when evaluated on class-balanced ID training data, they suffer from significant performance deterioration when the training sets follow a long-tailed distribution [42]. In such distributions, a few majority categories (head classes) dominate the dataset, while most minority categories (tail classes) have limited samples available. The primary challenges in long-tailed OOD detection arise from the confusion between OOD samples and head/tail classes, which manifests in two ways: (1) tail-class samples are erroneously treated as OOD samples [42, 8, 45, 28], and (2) OOD samples are often incorrectly classified into head classes [28]. Hence, it is imperative to effectively distinguish OOD samples from both head and tail classes. The separate class learning (SCL) approach has emerged as a promising solution, designed to scrutinize tail and head samples separately through tail-specific and head-specific class learning mechanisms. However, our analysis reveals two persistent limitations in SCL:
(a) Influence of static temperature
(b) Advantage of informative outlier mining
Figure 1: The OOD detection performance of representative methods (PASCL [42], TSCL [8]) and COCL [28] can be influenced by (a) static temperature values, and (b) informative outlier mining (OM) from the auxiliary OOD training data. The experiments are conducted on CIFAR100-LT benchmark. The AUROC and FPR95 are averaged over six OOD test sets (refer to Section 5.1).
1) Sensitive to static temperature: To distinguish OOD samples from head/tail classes, prevalent works predominantly rely on diverse augmentation within the realm of supervised contrastive learning [42, 8] or utilize learnable prototypes for tail classes [8, 28] to push head/tail samples away from OOD samples. Typically, the temperature parameter, which plays a pivotal role in regulating how much attention the model pays to samples, is employed as a static value across anchor, positive, and negative samples for simplicity. However, as illustrated in Figure 1(a), the OOD detection performance of these methods is influenced by varying temperature values. This sensitivity can be attributed to the use of a static temperature value, which results in uniform pulling and uniform pushing, a configuration that is incongruent with the intricacies of long-tailed data distribution. It is more reasonable to dynamically adjust the temperature value of each head and tail class to better align with the imbalanced distribution of class samples inherent in the training data.
2) Without consideration of inforamtive outliers: Recent works [42, 45, 28, 8] focus on improving long-tailed OOD detection performance through an approach called outlier exposure (OE) [12] that utilizes an unlabeled auxiliary training set as OOD training data. However, these methods have primarily relied on a random selection of outliers from the auxiliary OOD training data. This indiscriminate selection approach may lead to the inclusion of numerous uninformative outliers that lack significant alignment with either the head or tail classes. Such outliers exhibit restricted OOD characteristics, potentially yielding marginal enhancements to the model’s performance [2]. As illustrated in Figure 1(b), utilizing informative outliers mined from the OOD training data (refer to Section 4.3) can significantly improve the OOD detection performance. This highlights the advantages of enhancing the quality of auxiliary outliers.
To comprehensively address the aforementioned limitations of SCL, we introduce Adaptable Tail class Supervised Contrastive Learning (A-TSCL) and OOD-prototype-aware Head class Learning (A-OHL) as integral components of tail-specific and head-specific class learning, respectively. To mitigate these identified limitations, we propose Refined Separate Class learning (RSCL). Specifically, departing from the conventional static temperature value application across all ID classes, we propose dynamic class-wise temperature adjustment to modulate the temperature parameter for each head and tail class. This adaptive approach enables varying degrees of pulling within each tail class and distinct levels of pushing between OOD samples and head classes. In addition, as opposed to random outlier selection from the auxiliary OOD training set, we propose informative outlier mining to identify diverse types of outliers based on their affinity with head and tail classes. This will help to further enhance the efficacy of separate class learning. Through the above refinements to separate class learning, OOD samples are less prone to confusion with head and tail classes, thus improving the long-tailed OOD detection performance.
In conclusion, our main contributions are as follows:
  • We analyze the underlying limitations of the separate class learning approach for long-tailed OOD detection and reveal that the OOD detection performance is notably influenced by the presence of uninformative outliers and the use of static temperature parameters.
  • We propose refined separate class learning (RSCL), a novel approach for long-tailed OOD detection. RSCL utilizes dynamic class-wise temperature adjustment and informative outlier mining to mitigate the above limitations and effectively distinguish OOD samples from both head and tail classes.
  • Extensive evaluation and ablation studies demonstrate that RSCL achieves superior OOD detection performance while improving ID classification accuracy.
2 Related Work
OOD Detection.  Numerous methods have been developed to tackle OOD detection challenges within the realm of computer vision. One prevalent approach involves the utilization of post-hoc inference methods [11, 20, 24, 23, 14, 36, 41, 5, 50, 35, 22, 49]. These methods primarily focus on devising score functions that depend on features, logits, or softmax probabilities. Another approach involves training methods that do not rely on outlier data [13, 44, 29, 9, 37]. These methods incorporate training-time regularization techniques to enhance the OOD detection capabilities. The third approach entails training methods that utilize outlier data, assuming access to auxiliary OOD training samples [12, 24, 34, 30, 39, 2, 6, 38, 33, 43, 51, 26]. However, it is crucial to note that all these methods have primarily been developed and evaluated on datasets with balanced ID training data, thereby potentially exhibiting limitations when applied to imbalanced ID datasets.
Long-tailed OOD Detection.  Compared with OOD detection on balanced ID training data, significantly less work has been done on long-tailed scenarios. The pioneering work, PASCL [42], formulates the problem and reveals that simple combinations of existing OOD detection and long-tailed recognition methods [4, 15, 27] achieve unsatisfactory performance. PASCL identifies the challenge arising from the models’ struggle to differentiate between tail-class and OOD samples. To tackle this challenge, EAT [45] introduces dynamic virtual labels for OOD samples and augments the context-limited tail classes. TSCL [8] introduces an extra rejection class label for auxiliary OOD training data. COCL [28] recently demonstrates that OOD samples are often misclassified into head classes and proposes a novel approach called calibrated outlier class learning to distinguish OOD samples from both head and tail classes. However, these methods commonly adopt a random selection of outliers from the auxiliary OOD training data. We instead utilize informative outlier mining to identify and utilize diverse types of outliers based on their affinity with head and tail classes.
Supervised Contrastive Learning.  Supervised contrastive learning [17] has recently shown strong potential in enhancing long-tailed OOD detection. In its original formulation, all classes play the same role that pulling is done with every class and pushing between every pair of different classes. To distinguish OOD samples from tail classes, PASCL [42] proposes partial and asymmetric supervised contrastive learning, restricting contrastive interactions to tail classes and OOD samples. Building on this idea, TSCL [8] introduces tail-class prototype induced supervised contrastive learning to improve the performance of long-tail OOD detection further. Both PASCL and TSCL impose uniform pulling within each tail class and uniform pushing between OOD samples and tail classes, due to the employment of a static temperature value. In contrast, our approach employs dynamic class-wise temperature adjustment to modulate the temperature parameter for both head and tail class.
Figure 2: Overview of our approach RSCL. (a) High-level architecture of RSCL components. (b) Conceptual diagram of informative outlier mining: outliers are categorized as Tail-class-like OOD (black boundary), Neutral OOD (red boundary), and Head-class-like OOD (blue boundary). (c) Conceptual illustration of dynamic temperature adjustment for tail-specific class learning through A-TSCL (left diagram) and head-specific class learning via A-OHL (right diagram).
3 Preliminaries
Problem Statement.  Let $\mathcal{X}={X^{\text{in}}\cup X^{\text{out}}}$ represents the input space and $\mathcal{Y^{\text{in}}}=\left\{1,2,...,C\right\}$ be the set of $C$ imbalanced ID classes in the label space. The objective of long-tailed OOD detection is to learn a classifier $f$ that for any test data $\boldsymbol{x}\in\mathcal{X}$: if $\boldsymbol{x}$ is drawn from $X^{\text{in}}$ (from either head or tail classes), $f$ can classify $\boldsymbol{x}$ into the correct ID class; and if $\boldsymbol{x}$ is drawn from $X^{\text{out}}$, $f$ can identify $\boldsymbol{x}$ as OOD data. Typically, genuine OOD data is unavailable during the training phase, as OOD samples are unknown instances. However, auxiliary samples that do not belong to $X^{\text{out}}$ but are drawn from a different distribution other than $X^{\text{in}}$ are often accessible. These auxiliary samples can be utilized as pseudo-OOD samples to improve OOD detection performance and ID classification accuracy.
Outlier Class for Auxiliary OOD Training Data.  Outlier Exposure (OE) approach [12] is widely adopted for OOD detection on balanced ID datasets by utilizing an unlabeled auxiliary training set as OOD training data. However, when the training set follows a long-tailed distribution, the commonly adopted uniform distribution in OE becomes an ineffective prior. To address this issue, a straightforward solution is to learn an outlier class that encompasses the OOD samples [28, 8]. Specifically, for a classification problem with $C$ classes, the label space is extended by explicitly adding a separate class $(C+1)$ as an outlier class. During training, both the ID data $\mathcal{D}^{\text{train}}_{\text{in}}=(X_{\text{in}},Y_{\text{in}})$ and auxiliary data $\mathcal{D}^{\text{train}}_{\text{out}}=(X_{\text{out}},C+1)$ are utilized. The training objective is defined as follows:
$\mathcal{L}_{\text{OCL}}=\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}^{\mathrm{train}}_{\mathrm{in}}}[\ell(f(\boldsymbol{x}),y]+\alpha\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D}^{\mathrm{train}}_{\mathrm{out}}}[\ell(f(\boldsymbol{x}),\tilde{y}$], (1)
where $f(\boldsymbol{x})$ is the logits output, $\tilde{y}=C+1$, $\ell$ is a cross entropy loss, and $\alpha$ is a hyper-parameter.
4 Refined Separate Class Learning
We introduce a novel approach, termed Refined Separate Class Learning (RSCL), designed to mitigate the limitations in SCL. As shown in Figure 2(a), RSCL utilizes dynamic class-wise temperature adjustment and informative outlier mining to facilitate tail-specific and head-specific class learning. The components in RSCL are jointly trained with a shared feature extractor, ensuring comprehensive learning of representative features for all classes. Below we introduce each component in detail.
4.1 Separate Class Learning
Tail-specific Class Learning.  Due to the scarcity of tail-class samples in training data, models often mistakenly treat them as OOD samples [42, 45, 28, 8]. To mitigate this issue, an Adaptable Tail class Supervised Contrastive Learning (A-TSCL) is utilized to effectively distinguish tail classes from OOD samples. The formulation is as follows:
$\mathcal{L}_{\text{tail }}=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{I}}\left[\mathcal{L}_{\text{A-TSCL}}(\boldsymbol{x})\right],\text{ where }\mathcal{L}_{\text{A-TSCL}}(\boldsymbol{x})=$ (2)
[…]
where $\mathcal{P}(\boldsymbol{x})=\left\{\boldsymbol{p}\mid y(\boldsymbol{p})=y(\boldsymbol{x}),\boldsymbol{p}\in\mathcal{B}\backslash\{\boldsymbol{x}\}\right\}$ denotes a set of positive samples sharing the same ground-truth label as the tail sample $\boldsymbol{x}$, $\mathcal{B}$ is a batch of ID training samples from $\mathcal{D}^{\text{train}}_{\text{in}}$, $\mathcal{B}_{\text{t}}$ is a set of tail-class samples within $\mathcal{B}$, $\tilde{\mathcal{B}}$ is a batch of informative outliers from $\mathcal{D}^{\text{train}}_{\text{out}}$, $z(\boldsymbol{x})$ represents the output of a non-linear projection head on the model’s penultimate layer for sample $\boldsymbol{x}$, $\hat{\tau}_{c(\boldsymbol{x})}$ indicates a dynamic temperature parameter tailored for a tail class that is explicitly defined in Eq.(4), and $\mathcal{I}=\mathcal{B}_{\text{t}}\cup\mathcal{W}_{\text{t}}$ with $\mathcal{W}_{t}=(\boldsymbol{Z}_{t+1},\boldsymbol{Z}_{t+2},...,\boldsymbol{Z}_{C})\in\mathbb{R}^{(C-t)\times D}$ being a set of tail-class prototypes. Following TSCL [8], $\mathcal{W}_{\text{t}}$ is defined as trainable tail-class weight parameters of the model. To obtain $\mathcal{W}_{t}$, we first extract the tail-class weights $\mathcal{T}_{t}=(\boldsymbol{W}_{t+1},\boldsymbol{W}_{t+2},...,\boldsymbol{W}_{C})\in\mathbb{R}^{(C-t)\times D}$ from the classifier weights $\mathcal{W}=(\boldsymbol{W}_{1},\boldsymbol{W}_{2},...,\boldsymbol{W}_{C})\in\mathbb{R}^{C\times D}$. Subsequently, the tail-class weights in $\mathcal{T}_{t}$ are separately transformed by a multilayer perceptron (MLP). Finally, $\mathcal{W}_{t}$ are obtained after applying $\ell_{2}$ normalization.
The main difference between A-TSCL and previous works [42, 45, 28, 8] is the dynamic temperature parameter $\hat{\tau}_{c(\boldsymbol{x})}$, which is designed to adjust the temperature between tail samples during training.
Head-specific Class Learning.  Given the prevalent abundance of head-class samples in the training data, models often exhibit a prominent bias towards head classes when detecting OOD samples [28]. Consequently, OOD samples are frequently misclassified into head classes. Adaptable OOD-prototype-aware Head class Learning (A-OHL) is defined to effectively differentiate head classes from OOD samples. The formulation is as follows:
[…]
$\sum_{\tilde{\boldsymbol{x}}\in\tilde{\mathcal{B}}}-\frac{1}{|\tilde{\mathcal{B}}|}\log\frac{\exp(\frac{z(\tilde{\boldsymbol{x}})^{T}z(\tilde{\boldsymbol{w}})}{\tau})}{\sum\limits_{\boldsymbol{h}\in\mathcal{B}_{h}}\exp(\frac{z(\tilde{\boldsymbol{x}})^{T}z(\boldsymbol{h})}{\hat{\tau}_{c(\boldsymbol{h})}})+\exp(\frac{z(\tilde{\boldsymbol{x}})^{T}z(\tilde{\boldsymbol{w}})}{\tau})}.$
where $\tilde{\mathcal{B}}$ is a batch of informative outliers from $\mathcal{D}^{\text{train}}_{\text{out}}$, $\mathcal{B}_{\text{h}}$ is a set of head-class samples in $\mathcal{B}$, $\hat{\tau}_{c(\boldsymbol{h})}$ denotes a dynamic temperature parameter designed for a head class that is explicitly given in Eq.(4), and $\tilde{\boldsymbol{w}}$ represents the outlier-class prototype tailored for outliers. We define $\tilde{\boldsymbol{w}}$ as a trainable optimization parameter of the model and can be updated through gradient descent. Specifically, the weight associated with the outlier class (referred to as the (CC+1)-th class) is obtained from the (CC+1)-way classifier and then transformed through a projection head (MLP) to produce the outlier-class prototype, which enhancing the cohesion and compact representation of the outliers.
The main difference between A-OHL and previous works [42, 45, 28, 8] is the dynamic temperature parameter $\hat{\tau}_{c(\boldsymbol{x})}$, which is tailored for head classes and is utilized to adjust the temperature between a head-class sample and an OOD sample during training.
4.2 Dynamic Class-wise Temperature Adjustment
To fulfill the necessity of dynamic temperature adjustment in tail-specific and head-specific class learning in Section 4.1, we propose dynamic class-wise temperature adjustment, diverging from the conventional practice of employing a static temperature value across all sample pairs in previous works [42, 45, 28, 8]. The formulation is as follows:
$\hat{\tau}_{c(\boldsymbol{x})}=\tau\cdot\left[1-\sqrt{\frac{e}{E}}\cdot\sqrt{\hat{n}_{c(\boldsymbol{x})}}\right],(c(\boldsymbol{x})\in\{1,2,\ldots,C\})$ (4)
where $e$ denotes the current training epoch, $E$ indicates the total training epochs, $c(\boldsymbol{x})$ signifies the ground-truth label of sample $\boldsymbol{x}$, and $\tau$ is a static temperature value. We first construct a vector $\xi=\left(n_{1},n_{2},\ldots,n_{C}\right)$ based on the number of samples for each ID class, where $n_{i}(i\in\{1,2,\ldots,C\})$ represents the number of samples for the $i$-th class. For computational simplicity, we then subject $\xi$ to $\ell_{2}$ normalization to obtain $\hat{\xi}=\left(\hat{n}_{1},\hat{n}_{2},\ldots,\hat{n}_{C}\right)$.
Notably, our proposed temperature adjustment is contingent on the training epoch and the distributions of class samples. During training, the adjusted temperature value $\hat{\tau}$ gradually decreases, intensifying the pulling or pushing effect between the anchor and positive/negative sample. Specifically:
1) Regarding tail-specific class learning, as illustrated in left diagram in Fig.2(c), a tail-class sample is employed as an anchor, while other tail samples and tail-class prototypes assigned to each tail class are considered positive samples, and OOD samples are viewed as negative samples. Through the dynamic reduction of the temperature value $\hat{\tau}$ between the anchor and positive samples, we amplify the pulling effect within each tail class. This refinement augments the compactness of tail classes, rendering them more distinguishable from OOD samples.
2) For head-specific class learning, as illustrated in right diagram in Fig.2(c), an outlier serves as an anchor, with additional outliers and the outlier-class prototype acting as positive samples, while head-class samples are treated as negative samples. By dynamically decreasing the temperature value $\hat{\tau}$ between the anchor and negative samples, we intensify the pushing effect between OOD samples and head classes, thereby enhancing their distinguishability.
Through enhancing the feature representation of the tail class and reducing the overconfidence of the head class, the model can better distinguish OOD samples from both head and tail classes, thus improving the overall performance. Further discussion is provided in Appendix A.
4.3 Informative Outlier Mining
Instead of randomly sampling outliers from the auxiliary OOD training data [12, 24, 34, 30, 42, 45, 8, 28], we propose informative outlier mining to improve the quality of auxiliary outliers. The aim of informative outlier mining is to identify and utilize diverse types of outliers based on their affinity with head and tail classes. Each candidate outlier is evaluated through an outlier score function formulated as follows:
$\mathcal{S(\tilde{\boldsymbol{x}})}=1-\sum\nolimits_{n=1}^{N}\log\frac{\exp(f_{n}(\tilde{\boldsymbol{x}}))}{\sum_{j=1}^{C}\exp(f_{j}(\tilde{\boldsymbol{x}}))}$, (5)
[…]
In Eq.(5), we use the sum of softmax probabilities calculated over the ID tail classes as the outlier score, which serves as a measure of the likelihood of an outlier being classified as a tail or head class, or neither. Algorithm 1 in Appendix B provides the complete process of informative outlier mining.
As illustrated in Figure 2(b), after the outlier mining process, outliers are evenly classified into three distinct categories: 1) Tail-class-like OOD: Outliers that are closer in feature space to tail samples and are thus more likely to be misclassified as tail classes. 2) Head-class-like OOD: Outliers that exhibit higher similarity to head samples, making them prone to being predicted as head classes. 3) Neutral OOD: Outliers with no clear similarity to head or tail classes, equally likely to be misclassified as either. Further details can be found in Appendix B.
4.4 Training
The overall training objective of our approach is given as follows:
[…]
where $\mathcal{L}_{\text{OCL}}$ is defined in Eq.(1), $\mathcal{L}_{\text{tail}}$ represents the A-TSCL defined in Eq.(2), and $\mathcal{L}_{\text{head}}$ denotes the A-OHL defined Eq.(3). Hyper-parameters $\beta$ and $\gamma$ regulate the weighting of $\mathcal{L}_{\text{tail}}$ and $\mathcal{L}_{\text{head}}$.
The complete RSCL training process is detailed in Algorithm 2 in Appendix C. Specifically, during the first $\frac{3}{4}$ of the total training epochs, a subset of mixed outliers drawn from the three distinct OOD categories is utilized. By incorporating diverse outliers, the model is exposed to a broad spectrum of OOD features, thus facilitating the acquisition of comprehensive OOD knowledge. Subsequently, in the final $\frac{1}{4}$ of the total training epochs, outliers exclusively from the Neutral OOD category are utilized. These outliers are more ambiguous and challenging, and focusing on them in later training stages helps refine the model’s OOD discrimination capability. Further discussion on the advantage of utilizing diverse types of outliers at distinct training epochs is provided in Appendix E.1.
4.5 Inference
At inference time, a simple thresholding mechanism is used for detecting OOD samples. In particular, the OOD detector $G_{\eta}(\boldsymbol{x})$ can be constructed using the (CC+1)-th class as follows:
$G_{\eta}(\boldsymbol{x})=\begin{cases}1&\text{ if }F_{C+1}(\boldsymbol{x})\geq\eta\\ 0&\text{ if }F_{C+1}(\boldsymbol{x})<\eta,\end{cases}$ (7)
[…]
Table 1: OOD Detection Performance of RSCL, OE, and PASCL on CIFAR10/100-LT using ResNet-18. “Average” means the results averaged across six $\mathcal{D}_{\text{out}}^{\text{test}}$ sets. Bold numbers are the best results.
[…]
OE [12] / PASCL [42] / RSCL (Ours)
CIFAR10-LT
Texture 90.44 / 92.74 / 95.86 74.04 / 84.08 / 93.58 22.90 / 22.36 / 21.80
SVHN 92.39 / 95.54 / 98.43 93.27 / 97.37 / 99.35 16.14 / 14.01 / 07.59
CIFAR100 83.15 / 84.00 / 86.51 78.22 / 81.90 / 86.67 55.42 / 57.04 / 51.67
Tiny ImageNet 85.58 / 86.71 / 89.78 75.39 / 82.23 / 86.62 44.71 / 50.21 / 40.25
LSUN 90.88 / 92.93 / 95.99 85.71 / 91.17 / 95.70 26.00 / 26.69 / 18.85
Places365 89.25 / 91.17 / 94.19 93.95 / 96.02 / 97.76 31.70 / 31.99 / 26.25
Average 88.61 / 90.53 / 93.46 83.43 / 88.49 / 93.28 32.81 / 32.99 / 27.73
CIFAR100-LT
Texture 76.07 / 76.16 / 76.79 58.00 / 58.02 / 64.60 67.53 / 66.27 / 68.81
SVHN 78.53 / 78.72 / 85.31 87.42 / 87.12 / 90.93 58.19 / 54.78 / 42.46
CIFAR10 62.07 / 62.35 / 63.14 56.79 / 57.17 / 56.87 79.87 / 79.88 / 76.98
Tiny ImageNet 68.15 / 68.41 / 69.96 51.57 / 52.09 / 54.52 75.97 / 76.06 / 73.55
LSUN 76.80 / 77.11 / 81.78 60.50 / 61.47 / 68.49 64.38 / 63.72 / 55.62
Places365 75.49 / 75.85 / 79.68 86.10 / 86.53 / 89.15 65.30 / 65.50 / 59.73
Average 72.85 / 73.10 / 76.11 66.73 / 67.06 / 70.76 68.54 / 67.70 / 62.86
Table 2: Classification accuracy on head and tail classes of RSCL, OE, and PASCL.
Methods $\mathcal{D}_{\text{in}}^{\text{train}}$: CIFAR10-LT $\mathcal{D}_{\text{in}}^{\text{train}}$: CIFAR100-LT $\mathcal{D}_{\text{in}}^{\text{train}}$: ImageNet-LT
Head Tail Overall Head Tail Overall Head Tail Overall
OE 86.03 60.93 73.48 60.66 19.32 39.99 59.12 27.28 43.20
PASCL 86.44 65.44 75.94 60.55 19.35 39.95 60.98 29.51 45.29
RSCL(ours) 87.53 70.44 79.14 62.76 22.86 42.81 61.57 30.66 46.12
5 Experiments and Results
5.1 Experiment Settings
Datasets.  We use three popular long-tailed image classification datasets as ID training data (i.e., $\mathcal{D}_{\text{in}}^{\text{train}}$), including CIFAR10-LT, CIFAR100-LT [1], and ImageNet-LT [25]. For CIFAR10-LT and CIFAR100-LT, the imbalance ratio $\rho$ is set to 100 [27, 42]. The original CIFAR test sets and the ImageNet validation set serve as ID test data ($\mathcal{D}_{\text{in}}^{\text{test}}$). Following [12, 24, 42, 45], 300k Random Images is used as OOD training data ($\mathcal{D}_{\text{out}}^{\text{train}}$) for CIFAR10/100-LT and ImageNet-Extra [42] is used as $\mathcal{D}_{\text{out}}^{\text{train}}$ for ImageNet-LT. For OOD test set ($\mathcal{D}_{\text{out}}^{\text{test}}$), we use six datasets including Texture [3], SVHN [31], CIFAR [18], Tiny ImageNet [19], LSUN [48], and Places365 [52] introduced in the SC-OOD benchmark [46] for CIFAR10/100-LT. We use ImageNet-1k-OOD constructed by [42] as $\mathcal{D}_{\text{out}}^{\text{test}}$ to ImageNet-LT. More details about datasets are presented in Appendix D.1.
[…]
Implementation Details.  We compare our RSCL with several existing OOD detection methods on long-tailed training sets, including classical methods MSP [11], OE [12], EnergyOE [24], DAC [39], SOFL [30], NTOM [2], and very recently published methods PASCL [42], EAT [45], COCL [28], and TSCL [8]. For experiments on CIFAR10-LT and CIFAR100-LT, we use the standard ResNet18 [10] as the backbone. For experiments on ImageNet-LT, we follow the settings in [42, 45, 28] and use ResNet50 [10] as the backbone. Appendix D.2 provides more implementation details.
5.2 Main Results
[…]
Improvements on head and tail in-distribution classes.  Table 2 shows the classification accuracy of our RSCL over OE and PASCL on head and tail classes. Notably, our RSCL demonstrates substantial enhancements in classification accuracy across both head and tail classes, thereby benefiting the overall ID performance. Compared to PASCL[42], known for its pronounced bias towards tail classes with limited improvements in head classes, our RSCL achieves a good balance.
Comparison with other competitive methods.  Table 3 and Table 4 report the results for the CIFAR-LT and ImageNet-LT datasets, respectively. The results reveal that when the training set follows a long-tail distribution, existing post-hoc OOD detection methods [11] suffer a notable decline in their ability to detect OOD samples. Furthermore, methods that utilize auxiliary outlier data [39, 24, 30, 2] exhibit significant potential for improving OOD detection, but their performance on ID classification remains relatively poor. This is due to the absence of effective outlier mining that account for the affinity between outliers and both head and tail classes. Consequently, the model struggles to accurately differentiate between ID and OOD samples, resulting in misclassifications of ID samples. In particular, our RSCL also outperforms recently proposed methods [42, 45, 8, 28]. This consistent improvement on both ID and OOD data is due to our dynamic temperature adjustment and informative outlier mining, which enhances the discrimination ability of OOD samples from both head and tail classes.
Table 3: Comparison with other methods on CIFAR10-LT and CIFAR100-LT using ResNet18. Mean and standard deviation over six random runs are reported. Bold numbers are the best results.
Method $\mathcal{D}_{\text{in}}^{\text{train}}$: CIFAR10-LT $\mathcal{D}_{\text{in}}^{\text{train}}$: CIFAR100-LT
AUROC↑ AUPR↑ FPR95↓ ACC↑ AUROC↑ AUPR↑ FPR95↓ ACC↑
MSP [11] 68.68 67.69 75.20 69.84 61.38 58.22 83.15 41.14
EnergyOE [24] $87.98_{\pm 0.46}$ $84.55_{\pm 1.24}$ $37.71_{\pm 0.51}$ $66.53_{\pm 1.31}$ $68.76_{\pm 0.86}$ $62.15_{\pm 0.47}$ $72.87_{\pm 1.29}$ $33.87_{\pm 0.37}$
DAC [39] $89.91_{\pm 0.45}$ $86.26_{\pm 1.27}$ $31.99_{\pm 0.64}$ $75.30_{\pm 0.70}$ $72.51_{\pm 0.30}$ $66.60_{\pm 0.40}$ $68.96_{\pm 0.70}$ $41.54_{\pm 0.60}$
SOFL [30] $90.60_{\pm 0.56}$ $89.26_{\pm 1.44}$ $33.67_{\pm 0.86}$ $70.59_{\pm 1.14}$ $72.97_{\pm 0.78}$ $67.99_{\pm 1.08}$ $68.73_{\pm 0.80}$ $40.55_{\pm 0.84}$
NTOM [2] $90.61_{\pm 0.51}$ $90.19_{\pm 1.06}$ $36.53_{\pm 0.77}$ $68.78_{\pm 0.81}$ $70.90_{\pm 0.64}$ $66.76_{\pm 0.79}$ $72.59_{\pm 0.67}$ $35.68_{\pm 1.02}$
EAT [45] $90.98_{\pm 0.28}$ $89.34_{\pm 0.54}$ $33.01_{\pm 0.67}$ $78.74_{\pm 0.37}$ $72.97_{\pm 0.53}$ $67.75_{\pm 0.45}$ $67.91_{\pm 0.80}$ $42.71_{\pm 0.63}$
COCL [28] $91.66_{\pm 0.33}$ $91.06_{\pm 0.46}$ $34.40_{\pm 0.58}$ $74.87_{\pm 0.63}$ $74.68_{\pm 0.65}$ $\textbf{70.79}_{\pm 0.38}$ $67.87_{\pm 0.79}$ $41.80_{\pm 0.58}$
TSCL[8] $92.06_{\pm 0.08}$ $91.01_{\pm 0.25}$ $30.19_{\pm 0.36}$ $76.10_{\pm 0.51}$ $74.08_{\pm 0.30}$ $68.10_{\pm 0.21}$ $65.36_{\pm 0.68}$ $40.11_{\pm 0.26}$
RSCL(ours) $\textbf{93.46}_{\pm 0.06}$ $\textbf{93.28}_{\pm 0.09}$ $\textbf{27.73}_{\pm 0.25}$ $\textbf{79.14}_{\pm 0.34}$ $\textbf{76.11}_{\pm 0.34}$ $70.76_{\pm 0.33}$ $\textbf{62.86}_{\pm 0.58}$ $\textbf{42.81}_{\pm 0.40}$
Table 4: Results on ImageNet-LT using ResNet50. Bold numbers are the best results.
$\mathcal{D}_{\text{in}}^{\text{train}}$ Method AUROC↑ AUPR↑ FPR95↓ ACC↑
ImageNet-LT MSP[11] 53.81 51.63 90.15 39.65
OE [12] 68.36 70.46 87.31 43.20
EnergyOE [24] 75.57 74.28 78.18 35.67
DAC [39] 75.61 74.06 75.49 39.89
SOFL [30] 76.18 74.65 74.03 38.78
NTOM [2] - - - -
PASCL [42] 68.63 70.64 86.78 45.29
EAT [45] 69.55 69.10 86.66 43.74
COCL [28] 72.10 72.61 86.11 42.71
TSCL [8] 75.04 73.98 78.65 43.89
RSCL (ours) 78.90 78.10 73.04 46.19
Mitigating the confusion between OOD samples and head/tail-class samples.  Figure 3 presents a visualization of the feature embeddings corresponding to tail-class samples, head-class samples, and OOD samples. The feature distributions depicted in Figure 3(a) reveal that OOD samples are easily confused with some tail samples and a few head samples in long-tailed scenarios. However, when training the model with our RSCL objective in Eq.(6), a discernible enhancement in discriminating between OOD samples and head/tail samples is evident, as illustrated in Figure 3(b). This observation highlights the efficacy of our RSCL for long-tailed OOD detection.
(a) $\mathcal{L}_{\text{OCL}}$ Loss
(b) RSCL Loss (ours)
Figure 3: t-SNE [40] visualizations of penultimate-layer features from ResNet-18 trained on CIFAR10-LT using two different losses: (a) $\mathcal{L}_{\text{OCL}}$ loss as defined in Eq.(1), and (b) Our loss as defined in Eq.(6). We randomly select 300 head samples (75 per head class), 300 tail samples (50 per tail class), and 300 OOD samples from SVHN. The default value of $k$ is set to 0.6.
5.3 Ablation Studies
In this section, we provide the ablation results to analyze the influence of various factors in our approach. For consistency, our analysis is focused on the CIFAR100-LT benchmark. However, similar trends are observed on the CIFAR10-LT and ImageNet-LT benchmark as well.
Table 5: The effectiveness of different loss components. All values are percentages averaged over six OOD test datasets. $\mathcal{L}_{\text{tail}}$$^{\star}$ and $\mathcal{L}_{\text{head}}$$^{\star}$ represent the vanilla loss without incorporating the informative outlier mining (OM) and dynamic class-wise temperature adjustment (TA).
Loss Components AUROC↑ AUPR↑ FPR95↓ ACC↓
$\mathcal{L}_{\text{OCL}}$ $\mathcal{L}_{\text{tail}}$$^{\star}$ $\mathcal{L}_{\text{head}}$$^{\star}$ OM TA
$\checkmark$ 73.93 67.13 64.90 39.94
$\checkmark$ $\checkmark$ $\checkmark$ 74.15 68.34 65.81 40.11
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 75.29 69.25 63.33 41.70
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 75.23 70.29 64.88 42.05
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 74.87 68.09 63.63 40.25
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 75.06 69.72 64.57 43.02
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 76.11 70.76 62.86 42.81
The effectiveness of different loss components.  Our RSCL consists of three different loss components: (1) $\mathcal{L}_{\text{OCL}}$, (2) $\mathcal{L}_{\text{tail}}$ = ($\mathcal{L}_{\text{tail}}$$^{\star}$ + OM + TA), and (3) $\mathcal{L}_{\text{head}}$ = ($\mathcal{L}_{\text{head}}$$^{\star}$ + OM + TA). The ablation results of these three loss components are shown in Table 5. It is evident that the use of $\mathcal{L}_{\text{OCL}}$ in combination with either $\mathcal{L}_{\text{tail}}$ or $\mathcal{L}_{\text{head}}$ leads to marginal enhancements in OOD detection and ID classification. Optimal model performance is achieved when all three components are utilized simultaneously. This verifies the effectiveness of $\mathcal{L}_{\text{tail}}$ and $\mathcal{L}_{\text{head}}$ in improving model performance.
[…]
More Ablation Experiments. Additional ablation studies are provided in Appendix E.1.
6 Conclusion
This paper introduces Refined Separate Class Learning (RSCL), a novel approach designed to effectively distinguish OOD samples from both head and tail classes. RSCL addresses the constraints present in the original SCL approach by incorporating dynamic class-wise temperature adjustment and informative outlier mining. Specifically, dynamic class-wise temperature adjustment is proposed to modulate the temperature value for each head and tail class, while informative outlier mining is introduced to identify diverse types of outliers based on their affinity with head and tail classes. Comprehensive evaluation and ablation studies demonstrate that RSCL significantly enhances the OOD detection performance while improving the classification accuracy on in-distribution data.